A driver took 1.5 s to scan the instrument panel. During this time, the car was travelling at a speed of 105 km/h. The distance the car travelled while the driver scanned the panel was


73 m


49 m


70 m


44 m



1
Expert's answer
2021-11-08T16:49:08-0500
"d=vt=105\ \dfrac{km}{h}\times\dfrac{1000\ m}{1\ km}\times\dfrac{1\ h}{3600\ s}\times1.5\ s=44\ m."